Net Present Value Versus Internal Rate of Return For discount factors use Exhibit 12B-1 and Exhibit 12B-2. Skiba Company is thinking about two different modifications to its current manufacturing process. The after-tax cash flows associated with the two investments follow: Project I Project II $(100,000) 63,857 63,857 Year $(100,000) 134,560 Skibas cost of capital is 16%. Required 1. Compute the NPV and the IRR for each investment. Round present value calculations and your final NPV answers to the nearest dollar. Round IRR answers to the nearest whole percent. NPV IRR 161 % Project I Project II 2. Conceptual Connection: Explain why the project with the larger NPV is the correct choice for Skiba. NPV is an absolute profitabilitymeasure and reveals how much the value of the firm will change for each project. IRR gives a measure of relative profitability.Thus, since NPV reveals the total wealth change attributable to each project, it is preferred to the IRR measure. 13,793 4,766 X 18 %

Exhibit 12B.2 Present Value of an Annuity* 6 9% 10% 12% 14% 16% 18% 20% 25% 30% 2 1.97040 1.94156 1.91347 1.88609 1.85941 1.83339 1.80802 1.78326 1.759111.73554 1.69005 1.64666 1.60523 1.56564 1.52778 1.44000 1.36095 94099 2.88388 2.82861 2.77509 2.72325 2.67301 2.62432 2.57710 2.53129 2.48685 2.40183 2.32163 2.2458 5 4.85343 4.71346 4.57971 4.45182 4.32948 4.21236 4.10020 3.99271 3.88965 3.79079 3.60478 3.43308 3.27429 3.12717 2.99061 2.68928 2.43557 6 5.79548 5.60143 5.41719 5.24214 5.07569 4.91732 4.76654 4.62288 4.48592 4.35526 4.11141 3.88867 3.68474 3.49760 3.32551 2.95142 2.64275 76.72819 6.47199 6.23028 6.00205 5.78637 5.58238 5.38929 5.20637 5.03295 4.86842 4.56376 4.28830 4.03857 3.81153 3.60459 3.16114 2.80211 8 7.65168 7.32548 7.01969 6.73274 6.46321 6.20979 5.97130 5.74664 5.53482 5.33493 4.96764 4.63886 4.34359 4.077573.83716 3.32891 2.92470 56602 8.16224 7.78611 7.43533 7.10782 6.80169 6.51523 6.24689 32825 4.94637 4.60654 4.30302 4.03097 3.4631 3.01 12 11.25508 10.57534 9.95400 9.38507 8.86325 8.38384 7.9426 13 12.13374 11.34837 10.6349 8.30641 7.88687 7.49867 7.13896 6.80519 6.49506 5.93770 5.45273 5.02864 4.65601 4.32706 3.65640 3.14734 79322 4.43922 3.72512 3.19026 98565 9.39357 8.85268 8.35765 7.90378 7.48690 7.10336 6.42355 5.84236 5.34233 4.90951 4.53268 3.78010 3.22328 19711 15 13.86505 12.84926 11.93794 11.11839 10.37 16 14.71787 13.57771 12.56110 11.65230 10.83777 10.10590 9.44665 8.85137 8.31256 7.82371 6.97399 6.26506 5.66850 5.16235 4.72956 3.88741 3.28324 17 15.56225 14.29187 13.16612 12.16567 11.27407 10.47726 9.76322 9.12164 8.54363 8.02155 7.11963 6.37286 5.74870 5.22233 4.77463 3.90993 3.2948 141 19 17.22601 15.67846 14.32380 13.13394 12.08532 11.15812 10.33560 9.60360 8.95011 8.36492 7.36578 6.55037 20 18.04555 16.35143 14.87747 13.59033 12.46221 11.46992 10.59401 21 18.85698 17.01121 15.41502 14.02916 12.82115 11.76408 10.83553 10.01680 9.29224 8.64869 7.56200 6.68696 5.97314 5.38368 4.89132 3.96311 3.31984 5.93692 14.45112 13.16300 12.04158 11.06124 10.20074 944243 8.77154 7.6446 23 20.45582 18.29220 16.44361 14.85684 13.48857 12.30338 11.27219 10.37106 9.58021 8.88322 7.718 92453 3.97639 3.32535 514 62208 14.09394 12.78336 11.65358 10.67478 9.82258 9.07704 26 22.79520 20.12104 17.87684 15.98277 14.37519 13.00317 11.82578 10.80 27 23.55961 20.70690 18.32703 16.32959 14.64303 13.21053 11.98671 10.93516 10.02658 9.23722 7.94255 28 24.31644 21.28127 18.76411 16.66306 14.89813 13.40616 12.13711 11.05108 10.11 29 25.06579 21.84438 19.18845 16.98371 15.14107 13.59072 12.27767 11.15841 10.19828 9.36961 8.0218 30 25.80771 22.39646 19.60044 17.29203 15.37245 13.76483 12.40904 11.25778 10.27365 9.42691 8.05518 7.00266 6.17720 5.51681 4.97894 3.99505 3.33206 8304 6.16555 5.50983 4.97472 3.99381

Exhibit 12B.1 Present Value of a Single Amount* 9 25° 0.96154 0.95238 0.94340 0.93458 0.92593 0.91743 0.90909 0.8 344 62 0.81630 0.79383 0.77218 0.75131 0.71178 0.67497 0.64066 0.60863 0.57870 0.51200 0.45517 63552 0.59208 0.55229 0.51579 0.48225 0.40960 0.35013 0.85480 0.82270 0.79209 0.76290 0.73503 0.70843 0.68301 0.94205 0.88797 0.83748 0.79031 0.74622 0.7 0.66634 0.63017 0.59627 0.56447 0.50663 0.45559 0.41044 0.37043 0.33490 0.26214 0.2071 0.66506 0.62275 0.58349 0.54703 0.51316 0.45235 0.39964 0.35383 0.31393 0.27908 0.20972 0.15937 30503 0.26604 0.23257 0.16777 0.12259 8 0.92348 0.85349 0.78941 0.73069 0.67684 0.62741 0.58201 0.54027 0.50187 0.46651 0.40388 0.3505 7025 67556 0.61391 0.55839 0.50835 0.46319 0.42241 0.38554 0.32197 0.26974 0.226 0.19106 0.16151 0.10737 0.07254 11 0.89632 0.80426 0.72242 0.64958 0.58468 0.52679 0.47509 0.42888 0.38753 0.35049 0.28748 0.23662 0.19542 0.16192 0.13459 0.08590 0.0558 16846 0.13722 0.11216 0.06872 0.04292 0.66112 0.57748 0.50507 0.44230 0.38782 0.34046 0.29925 0.26333 0.20462 0.15971 0.12520 0.09 4301 0.64186 0.55526 0.48102 0.41727 0.36245 1524 0.27454 0.23939 0.18270 0.14010 0.10793 0.08352 0.06491 0.03518 8 2 4438 0.71416 0.60502 0.51337 0.43630 0.37136 0.31657 0.27027 0.23107 0.19784 0.14564 0.10780 0.08021 0.059 83602 0.70016 0.58739 0.49363 0.41552 0.35034 0.29586 0.25025 0.21199 0.17986 0.13004 456 0.06914 0.05083 0.03756 0.01801 0.00889 954 0.67297 0.55368 0.45639 0.376 0.31180 0.25842 0.21455 0.17843 0.14864 0.10367 0.07276 0.05139 0.03651 0.02608 0.01153 0.0052 894 0.29416 0.24151 0.19866 0.16370 0.13513 0.09256 0.06383 0.04430 0.03094 0.02174 0.00922 0.00405 22 0.80340 0.64684 0.52189 0.42196 0.34185 0.27751 0.22571 0.18394 0.15018 0.12285 0.08264 0.05599 φ0819 0.02522 00131, 000738 0.003m 5 24 0.78757 0.62172 0.49193 0.39012 0.31007 0.24 25 0.77977 0.60953 0.47761 0.37512 0.29530 0.23300 0.18425 0.14602 0.11597 0.09230 0.05882 0.03779 0.02447 0.01596 0.01048 0.00378 0.00142 19715 0.15770 0.12640 0.10153 0.06588 0.04308 0.02838 0.01883 0.01258 0.00472 0.00184 24 0.21981 0.17220 0.13520 0.10639 0.08391 0.05252 0.03315 0.0210 27 0.76440 0.58586 0.45019 0.34682 0.26785 0.20737 0.16093 0.12519 0.09761 0.07628 0.04689 0.02908 0.01818 0.01146 0.00728 0.00242 0.000 28 0.75684 0.57437 0.43708 0.33348 0.25509 0.19563 0.15040 0.11591 0.08 4 0.06934 0.04187 0.02551 0.01567 0.00971 0.00607 0.00193 0.00065 11 0.42435 0.32065 0.24295 0.18456 0.14056 0.10733 0.08215 0.06304 0.03738 0.02237 0.01351 0.00823 0.00506 0.00155 0.00050 3 0.
Answer #1

Solution to 1- Net Present Value method of capital budgeting uses the discount rates specified to calculate the future values of cash inflows to present discounted value at the time of actual cash outflows.

Calculation of Net Present Value:

NPV = Present value of cash inflows - Present value of cash outflows

(Present value is calculated with the help of Exhibit 12B.1)

Project 1- Present value of cash inflows = 134560*0.74316 = 99999.6 (Rounding of to nearest dollar=100000)

Cash outflows = 100000

NPV = 100000 - 100000 =0

Therefore, NPV of project 1 is 0

Project 2- Present value of cash inflows = (63857*0.86207) + (63857*0.74316) = 102505.16 (Rounding of=102505)

Cash outflows = 100000

NPV = 102505 - 100000 = 2505

Calculation of Internal Rate of Return:

Internal Rate of Return is a discount rate that equates the present value of expected cash outfows with cash inflows

Therefore,Internal Rate of Return = cash inflows/cash outflows = 1

(IRR is calculated by basic trial and error by the help of Exhibit 12B.1)

Project 1- Present value of cash inflows =134560*0.74316 = 100000 (99999.6) (16% is taken by trial and error as it was closest to equate present value of cash inflows with cash outflows)

Cash outflows = 100000

Therefore Internal rate of return for project 1 = 16%

Project 2- Present value of cash inflows = (63857*0.86207) + (63857*0.74316) = 102505 (16% is substituted by trial and error)

But as we can see that 16% is not equating present value of cash inflows with cash outflows,therefore we will substitute 18% is it will lower the value of cash inflows

Present value of cash inflows = (63857*0.84746) + (63857*0.71818) = 99980 (18% is substituted by trial and error)

Internal rate of return = 16% + (102505-100000)/102505-99980)*(18-16) = 16% + 1.984% = 17.984%(Rounding of to nearest percentage = 18%)

Therefore Internal rate of return for project 2 = 18%

Solution to 2-  NPV and IRR differ in the sense that result under certain circumstances are mutually contradictory under two methods.This happens mostly in the case of mutually exclusive proposals,where one favours one project and the other favours the other project. NPV is an gives an absolute result in the form of cash flows whereas IRR provides a relative result in the form of percentages.therefore the decision making become easier as all we have to do is to prefer the project with higher NPV.